OpenStudy (datboydev):

Determine the rate law, including the values of the orders and rate law constant, for the following reaction using the experimental data provided. X + Y yields products

OpenStudy (cuanchi):

you haven't provide the experimental data. Usually is a table like this experiment [X], M [Y], M rate = -d[X]/dt, M h-1 1 0.5 0.5 1.2 2 1 0.5 4.8 3 2 1 38.4

OpenStudy (datboydev):

sorry ill put that now'

OpenStudy (datboydev):

Trial [X] [Y] Rate 1 0.20 M 0.15 M 2.4 × 10-2 M/min 2 0.20 M 0.30 M 4.8 × 10-2 M/min 3 0.40 M 0.30 M 19.2 × 10-2 M/min

OpenStudy (datboydev):

@cuanchi

OpenStudy (cuanchi):

take the two first esperiments Trial [X] [Y] Rate 1 0.20 M 0.15 M 2.4 × 10-2 M/min 2 0.20 M 0.30 M 4.8 × 10-2 M/min the concentration of X is the same and the concentration of Y increase twice, the rate also increase twice then the order for Y will be =1 Rate = k [X]^m [Y]^n n= 1 then you take the Trials 2 and 3 2 0.20 M 0.30 M 4.8 × 10-2 M/min 3 0.40 M 0.30 M 19.2 × 10-2 M/min here the concentration od Y is constant and the concentration of X change to double and the rate increase 4 times then the order for X is....?????

OpenStudy (datboydev):

i don't really understand this lesson

OpenStudy (datboydev):

so is it 2

OpenStudy (datboydev):

m=2???

OpenStudy (datboydev):

@cuanchi what do i do after since i have n=1 m=2?

OpenStudy (cuanchi):

write the rate law Rate = k [X]^m [Y]^n replace the m and n for the values

OpenStudy (cuanchi):

also you can calculate the value of k

OpenStudy (datboydev):

so k[A]^2[B]^1

OpenStudy (datboydev):

then i use one of the trials above right

OpenStudy (cuanchi):

yes

OpenStudy (datboydev):

so [.20]^2[.15]^2

OpenStudy (datboydev):

[.15]^1

OpenStudy (datboydev):

2.4 × 10-2 M/min=k[.20]^2[.15]^1

OpenStudy (datboydev):

what do i do now? @cuanchi

OpenStudy (datboydev):

k=2.4 ?/m^3s^3

OpenStudy (datboydev):

i mean M^3 m^3

OpenStudy (datboydev):

i think i got it is it k=2.4 4/M^3 m^3

OpenStudy (datboydev):

@sweetburger

OpenStudy (cuanchi):

isolate the k

OpenStudy (datboydev):

i dont know what to do after that to be honest

OpenStudy (cuanchi):

k= Rate / ([X]^m [Y]^n)

OpenStudy (datboydev):

so its not 1?

OpenStudy (sweetburger):

Oh i was looking at this " [.20]^2[.15]^2"

OpenStudy (sweetburger):

If you determined its 1. Then you are correct.

OpenStudy (datboydev):

ok

OpenStudy (datboydev):

k= 2.4 × 10-2 M/min 4/M^3 m^2?

OpenStudy (cuanchi):

NO

OpenStudy (cuanchi):

k= Rate / ([X]^m [Y]^n) k= 2.4 × 10-2 M/min / ((0.20 M ^1) (0.15 M ^2))

OpenStudy (cuanchi):

\[k =\frac{ 2.4 \times 10^{-2}M/\min }{ (0.2M)^{1}\times(0.15M)^{2} } \]

OpenStudy (datboydev):

k=.0027

OpenStudy (cuanchi):

to double check you can replace this value in the rate law and calculate the rate with the concentrations of the Trial 2 Rate = k [X]^m [Y]^n 2 0.20 M 0.30 M 4.8 × 10-2 M/min and see if you get 4.8 × 10-2 M/min

OpenStudy (datboydev):

same thing

OpenStudy (datboydev):

thx